OpenStudy (anonymous):

What is the percent of increase from 25 to 35? A. 10% B. 29% C. 40% D. 71%

OpenStudy (anonymous):

Percentage: initial/final Thus (25/35)100%

OpenStudy (anonymous):

is it A?

OpenStudy (anonymous):

What is your reasoning behind choosing A?

OpenStudy (anonymous):

well all i did was add 10 to 25

OpenStudy (anonymous):

is it the right answer?

OpenStudy (anonymous):

That doesn't give you a percentage, that just gives the difference between the two values. If you take 10% of 35, (35*.1), you get 3.5 - therefore 10% is not the right answer. Try using the formula I already gave you

OpenStudy (anonymous):

ohh its C

OpenStudy (anonymous):

What reasoning gave you C?

OpenStudy (anonymous):

p = (35 - 25)/25 p = 10/25 = 2/5 p = 0.4 p = 40%

OpenStudy (anonymous):

There ya go

OpenStudy (anonymous):

thanks for the help